OpenStudy (anonymous):

Describe the differences between the graph of y equals negative 3 over 4 open parentheses x minus 6 close parentheses squared plus 4 and the standard position graph of y = x2. What are the x- and y-intercepts of the quadratic equation y = x2 + 2x - 3? What are the x- and y-intercepts of the quadratic equation y = (x - 1)2 - 16? Identify the x-intercepts by graphing the quadratic equation.

OpenStudy (anonymous):

What are the x- and y-intercepts of the quadratic equation y = x2 + 2x - 3? (x-1)(x-2)=y So, x intercepts are (0,1), (0,2) Y is (2,0)

OpenStudy (anonymous):

I lied, x intercepts are (1,0), (2,0). Y is (0,2)

OpenStudy (anonymous):

What are the x- and y-intercepts of the quadratic equation y = (x - 1)2 - 16? equals x^2 - 2x - 15 (x+5)(x-3) x intercepts are (-5,0),(3,0) y is at (0, -15)